# Relax and Merge: A Simple Yet Effective Framework for Solving Fair $k$-Means and $k$-sparse Wasserstein Barycenter Problems

### v1

@@ 1 Introduction
Clustering is one of the most fundamental problems in the area of machine learning. A wide range of practical applications rely on effective clustering algorithms, such as feature engineering (Glassman et al., 2014; Alelyani et al., 2018), image processing (Coleman and Andrews, 1979; Chang et al., 2017), and bioinformatics (Ronan et al., 2016; Nugent and Meila, 2010). In particular, the $k$-means clustering problem has been extensively studied in the past decades (Jain, 2010). Given an input dataset $P\subset\mathbb{R}^{d}$, the goal of the $k$-means problem is to find a set $S$ of at most $k$ points for minimizing the clustering cost, which is the sum of the squared distances from every point of $P$ to its nearest neighbor in $S$. In recent years, motivated by various fields like education, social security, and cultural communication, the study on fairness of clustering has in particular attracted a great amount of attention (Chierichetti et al., 2017; Bera et al., 2019; Huang et al., 2019; Chen et al., 2019; Ghadiri et al., 2021).
@@
In this paper, we consider the problem of $(\alpha,\beta)$-fair $k$-means clustering that was initially proposed by Chierichetti et al. (2017) and then generalized by Bera et al. (2019). Informally speaking, we assume that the given dataset $P$ consists of $m$ groups of points, and the “fairness” constraint requires that in each obtained cluster, the points from each group should take a fraction between pre-specified lower and upper bounds. Bera et al. (2019) showed that a $\rho$-approximate algorithm for vanilla $k$-means can provide a $(2+\sqrt{\rho})^{2}$- approximate solution  [^1] In their paper, the approximate ratio is written as $(2+\rho)$ because they added a squared root to the $k$-means cost function. for $(\alpha,\beta)$-fair $k$-clustering with a slight violation on the fairness constraints, where the “violation” is formally defined in Section 2. Furthermore, Böhm et al. (2021) studied the “strictly” fair $k$-means clustering problem, where it requires that the number of points from each group should be uniform in every cluster; they obtained a $(2+\sqrt{\rho})^{2}$ approximate solution without violation. These fair $k$-means algorithms can also be accelerated by using the coreset techniques, such as (Huang et al., 2019; Braverman et al., 2022; Bandyapadhyay et al., 2024). There also exist polynomial-time approximation scheme (PTAS) for fair $k$-means, such as the algorithms proposed in (Böhm et al., 2021; Schmidt et al., 2020; Bandyapadhyay et al., 2024), but their methods have an exponential time complexity in $k$. We are also aware of several other different definitions of fairness for clustering problems, such as the proportionally fair clustering (Chen et al., 2019; Micha and Shah, 2020) and socially fair $k$-means clustering (Ghadiri et al., 2021).
@@ 4 Experiments
In this section, we perform the empirical evaluation on our algorithms. Our experiments are conducted on a server equipped with Intel(R) Xeon(R) Gold 6154 CPU @ 3.00GHz CPU and 512GB memory. We implement our algorithms in C++ and python (with linear programming solver gurobi (Gurobi Optimization, LLC, 2023)). We use the following datasets which are commonly used in previous works: Bank (Moro et al., 2014)(4522 points with 5 groups), Adult (Becker and Kohavi, 1996) (32561 points with 7 groups), Census (Zhou and Chen, 2002)(50000 points with 10 groups), creditcard (Yeh and Lien, 2009) (30000 points with 8 groups), Biodeg (Mansouri et al., 2013) (1055 points with 2 groups), Breastcancer (Wolberg,William, Mangasarian,Olvi, Street,Nick, and Street,W., 1995) (570 points with 2 groups), Moons (scikit-learn developers, 2007-2023) (200 points with 2 groups), Hypercube(200 points with 2 groups), Cluto (Karypis et al., 1999) (800 points with 8 groups), and Complex (800 points with 8 groups). The last four datasets consist of disjoint and equal sized groups, so we can perform strictly fair $k$-means algorithms on them. We place the detailed information of these datasets in Appendix E. Regarding the selection of $\alpha$ and $\beta$, we set $\alpha_{i}=\beta_{i}=\frac{|P^{(i)}|}{|P|}$ and we also discuss more choices for $\alpha$ and $\beta$, and provide more experimental results, including the part of $k$-sparse Wasserstein Barycenter, in the Section E of the appendix.
@@ Appendix D Fixed Support Wasserstein Barycenter
Appendix E Supplementary Experiment
E.1 Datasets
@@
E.2 Comparison on Cost with different $k$ and $(\alpha,\beta)$
@@
E.3 Comparison on Cost of $k$-sparse Wasserstein Barycenter
@@
E.4 Cost on Different sampling ratio
@@
E.5 Running time with different sampling ratio on $T$
As we discussed in E.4, sampling on relaxed solution $T$ can reduce the running time while the overall cost not increasing too much. We also test the running time with different sampling ratio. In summary, the running time of solving LP(2) on $T$ and overall Algorithm 1, shown in Table 3 and Table 4, can be significantly reduced by sampling.
@@ E.5 Running time with different sampling ratio on $T$
E.6 Experiments of Our Rounding Algorithm
@@
In this section, we implement our rounding algorithm in Appendix C and compute the violation factor across different datasets and parameters. For convenience, we parameterize $\alpha_{i}$ and $\beta_{i}$ for the $i$-th group using a single parameter $\delta$. Specifically, we set $\beta_{i}=\frac{|P^{(i)}|(1-\delta)}{|P|}$ and $\alpha_{i}=\frac{|P^{(i)}|}{|P|(1-\delta)}$. Generally speaking, the smaller the $\delta$, the stricter the fairness constrains are. In Table 5 6 7, the violation introduced by our rounding algorithm is less than 1 in most of the cases and never exceeds 2, which aligns with our theoretical analysis.

### v3

@@ 1 Introduction
Clustering is one of the most fundamental problems in the area of machine learning. A wide range of practical applications rely on effective clustering algorithms, such as feature engineering (Glassman et al., 2014; Alelyani et al., 2018), image processing (Coleman and Andrews, 1979; Chang et al., 2017), and bioinformatics (Ronan et al., 2016; Nugent and Meila, 2010). In particular, the $k$-means clustering problem has been extensively studied in the past decades (Jain, 2010). Given an input dataset $P\subset\mathbb{R}^{d}$, the goal of the $k$-means problem is to find a set $S$ of at most $k$ points for minimizing the clustering cost, which is the sum of the squared distances from every point of $P$ to its nearest neighbor in $S$. In recent years, motivated by various fields like education, social security, and cultural communication, the study on fairness of clustering has in particular attracted a great amount of attention (Chierichetti et al., 2017; Bera et al., 2019; Huang et al., 2019; Chen et al., 2019; Ghadiri et al., 2021; Backurs et al., 2019).
@@
In this paper, we consider the problem of $(\alpha,\beta)$-fair $k$-means clustering that was initially proposed by Chierichetti et al. (2017) and then generalized by Bera et al. (2019). Informally speaking, we assume that the given dataset $P$ consists of $m$ groups of points, and the “fairness” constraint requires that in each obtained cluster, the points from each group should take a fraction between pre-specified lower and upper bounds. Bera et al. (2019) showed that a $\rho$-approximate algorithm for vanilla $k$-means can provide a $(2+\sqrt{\rho})^{2}$- approximate solution  [^1] In their paper, the approximate ratio is written as $(2+\rho)$ because they added a squared root to the $k$-means cost function. for $(\alpha,\beta)$-fair $k$-clustering with a slight violation on the fairness constraints, where the “violation” is formally defined in Section 2. Regarding to the no violation scenario, Dai et al. (2022) and Wu et al. (2024) both obtained an $O(\log k)$-approximate solution for fair $k$-median, and their basic technique is dynamic programming (Arora et al., 1998) after embedding the metric space into a tree. Wu et al. (2024) also achieved a quasi-polynomial-time approximate scheme in doubling metric by using dynamic programming with split tree (Talwar, 2004). Furthermore, Böhm et al. (2021) studied the “strictly” fair $k$-means clustering problem, where it requires that the number of points from each group should be uniform in every cluster; they obtained a $(2+\sqrt{\rho})^{2}$ approximate solution without violation. These fair $k$-means algorithms can also be accelerated by using the coreset techniques, such as (Huang et al., 2019; Braverman et al., 2022; Bandyapadhyay et al., 2024). There also exist polynomial-time approximation scheme (PTAS) for fair $k$-means, such as the algorithms proposed in (Böhm et al., 2021; Schmidt et al., 2020; Bandyapadhyay et al., 2024), but their methods have an exponential time complexity in $k$. We are also aware of several other different definitions of fairness for clustering problems, such as the proportionally fair clustering (Chen et al., 2019; Micha and Shah, 2020) and socially fair $k$-clustering clustering (Ghadiri et al., 2021; Abbasi et al., 2021; Makarychev and Vakilian, 2021; Chlamtáč et al., 2022).
@@ 4 Experiments
In this section, we perform the empirical evaluation on our algorithms. Our experiments are conducted on a server equipped with Intel(R) Xeon(R) Gold 6154 CPU @ 3.00GHz CPU and 512GB memory. We implement our algorithms in C++ and python (with linear programming solver gurobi (Gurobi Optimization, LLC, 2023)). We use the following datasets which are commonly used in previous works: Bank (Moro et al., 2014)(4522 points with 5 groups), Adult (Becker and Kohavi, 1996) (32561 points with 7 groups), Census (Zhou and Chen, 2002)(50000 points with 10 groups), creditcard (Yeh and Lien, 2009) (30000 points with 8 groups), Biodeg (Mansouri et al., 2013) (1055 points with 2 groups), Breastcancer (Wolberg,William, Mangasarian,Olvi, Street,Nick, and Street,W., 1995) (570 points with 2 groups), Moons (scikit-learn developers, 2007-2023) (200 points with 2 groups), Hypercube(200 points with 2 groups), Cluto (Karypis et al., 1999) (800 points with 8 groups), and Complex (800 points with 8 groups). The last four datasets consist of disjoint and equal sized groups, so we can perform strictly fair $k$-means algorithms on them. We place the detailed information of these datasets in Appendix F. Regarding the selection of $\alpha$ and $\beta$, we set $\alpha_{i}=\beta_{i}=\frac{|P^{(i)}|}{|P|}$ and we also discuss more choices for $\alpha$ and $\beta$, and provide more experimental results, including the part of $k$-sparse Wasserstein Barycenter, in the Section F of the appendix. We use $k$-means++ (Ostrovsky et al., 2013) as the $k$-means solver in our Algorithm 1.
@@ Appendix D Fixed Support Wasserstein Barycenter
For the sake of completeness, we need to clarify how the solution to the $k$-sparse Wasserstein barycenter solution is guaranteed to be a distribution. After we run Algorithm 1, we obtain the support $S$ (the locations of centers) of the returned solution and the assignment matrix $\phi_{S}^{*}$ (the transportation weight from $p=P^{(l)}_{i}$ to $f=S_{j}$ is denoted by $\phi^{*}_{S}(p,f)=x^{(l)}_{ij}$ in (16)). The key question is how to ensure that the summation of the weight of points in $S$ is equal to 1. Let us consider an arbitrary given distribution (or ”group” in the context of fair $k$-means), e.g., $P^{(l)}$. For every facility $f$ in $S$, we define its weight $w(f)=\sum_{p\in P^{(l)}}\phi^{*}_{S}(p,f)$. This ensures that the total weight of $S$ must be equal to the total weight of $P^{(l)}$, which is 1 because $P^{(l)}$ is a distribution. The choice of $P^{(l)}$ can be arbitrary because, recall that $k$-sparse WB can be seen as a special fractional version of strictly fair $k$-means, meaning no matter which given distribution you choose, you will obtain the same weight distribution of $S$. The optimization will not change by setting the weight of $S$ because the weight of $S$ does not affect the cost.
Appendix E Extend Algorithm 1 to $k$-Median and $k$-Means in General Metric Space
Although we mainly consider the fair $k$-means problem in Euclidean space in this paper, for the sake of completeness, in this section, we illustrate how to extend our framework to solve $k$-median and $k$-means in general metric space. In summary, if the potential facility set is given, our framework achieves a $(1+2\rho)$-approximate solution for $k$-median ($(2+8\rho)$-approximate solution for $k$-means) in metric space, where $\rho$ is the approximation ratio for vanilla $k$-median ($k$-means) with a constant violation factor. If the metric space has a fixed doubling dimension [Gupta et al., 2003], then equipped with existing PTAS for metric $k$-median and $k$-means [Cohen-Addad et al., 2021, 2019, Friggstad et al., 2019], the best approximation ratios our framework can achieve are $(3+O(\epsilon))$ for fair $k$-median and $(10+O(\epsilon))$ for fair $k$-means.
Unfortunately, our theoretical guarantees in general metric space are weaker than those of Bera et al. [2019], in which they obtained a $(\rho+2)$-approximation for $k$-median and a $(\sqrt{\rho}+2)^{2}$-approximation for $k$-means. The obstacle to achieving a better approximation ratio for our framework is the ”candidate set”. In Euclidean space, we have an approximate centroid set. However, in general metric space, how can we obtain a candidate set that has similar properties to Proposition 1, which provides a more powerful tool than the basic triangle inequality? This is not only a potential future work of our framework but also an important open theoretical problem.
$k$-Median in metric space.
Firstly, we consider fair $k$-median in general metic space. We use $\mathtt{dist}(\cdot,\cdot)$ to denote the distance between two points. We assume that the potential facility set $T$ is given. Therefore, in Algorithm 1, we just use the given facility set $T$ rather than computing the approximate centroid set. The cost of fair $k$-median can be written as
$\mathtt{Cost}(P,S,\phi_{S}^{*})$ $=\sum_{p\in P}\sum_{s\in S}\mathtt{dist}(p,s)\phi_{S}^{*}(p,s).$ (17)
Similar to Lemma 1, we have the following lemma.
Lemma 7.
Let $\eta$ be any positive number. If we suppose $\mathtt{Cost}(P,T,\phi_{T}^{*})\leq\eta\cdot OPT$, then the solution $(S,\phi_{S}^{*})$ returned by Algorithm 1 (the construction of $T$ should be slightly changed) is an $\big{(}\eta+(\eta+1)\rho\big{)}$-approximate solution for fair $k$-median problem in metric space, where $\rho$ is the approximation ratio of vanilla $k$-median.
Proof.
$\mathtt{Cost}(P,S,\phi_{S}^{*})$ $\leq\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(p,\mathcal{N}(t,S))\phi_{T}^{*}(p,t)$ (18)
$\leq\sum_{t\in T}\sum_{p\in P}\Big{[}\mathtt{dist}(p,t)+\mathtt{dist}(t,\mathcal{N}(t,S))\Big{]}\phi_{T}^{*}(p,t)$
$=\underbrace{\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(p,t)\phi_{T}^{*}(p,t)}_{\text{(a)}}+\underbrace{\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(t,\mathcal{N}(t,S))\phi_{T}^{*}(p,t)}_{(b)}.$
The second inequality is triangle inequality. Then we bound (a) and (b) separately. Firstly,
$(a)=\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(p,t)\phi_{T}^{*}(p,t)=\mathtt{Cost}(P,T,\phi^{*}_{T})\leq\eta\cdot OPT$ (19)
Next, we focus on (b). Suppose $S_{median}$ is the optimal $k$-median solution of $T$. Then we have:
$(b)$ $=\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(t,\mathcal{N}(t,S))\phi_{T}^{*}(p,t)$ (20)
$\leq\rho\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(t,\mathcal{N}(t,S_{median})\phi_{T}^{*}(p,t)$
$=\rho\sum_{p\in P}\sum_{t\in T}\big{[}\sum_{\tilde{s}\in S_{opt}}\mathtt{dist}(t,\mathcal{N}(t,S_{median}))\phi_{S_{opt}}^{*}(p,\tilde{s})\big{]}\phi_{T}^{*}(p,t)$
$\leq\rho\sum_{p\in P}\sum_{t\in T}\big{[}\sum_{\tilde{s}\in S_{opt}}\mathtt{dist}(t,\tilde{s})\phi_{S_{opt}}^{*}(p,\tilde{s})\big{]}\phi_{T}^{*}(p,t).$
Further, according to the triangle inequality, we have
$(b)$ $\leq\rho\sum_{p\in P}\sum_{t\in T}\Big{[}\sum_{\tilde{s}\in S_{opt}}\big{[}\mathtt{dist}(t,p)+\mathtt{dist}(p,\tilde{s})\big{]}\phi_{S_{opt}}^{*}(p,\tilde{s})\Big{]}\phi_{T}^{*}(p,t)$ (21)
$\leq\rho\sum_{p\in P}\sum_{t\in T}\sum_{\tilde{s}\in S_{opt}}\mathtt{dist}(t,p)\phi_{S_{opt}}^{*}(p,\tilde{s})\phi_{T}^{*}(p,t)$
$+\rho\sum_{p\in P}\sum_{t\in T}\sum_{\tilde{s}\in S_{opt}}\mathtt{dist}(p,\tilde{s})\phi_{S_{opt}}^{*}(p,\tilde{s})\phi_{T}^{*}(p,t)$
$=\rho\sum_{p\in P}\sum_{t\in T}\mathtt{dist}(t,p)\phi_{T}^{*}(p,t)+\rho\sum_{p\in P}\sum_{\tilde{s}\in S_{opt}}\mathtt{dist}(p,\tilde{s})\phi_{S_{opt}}^{*}(p,\tilde{s}).$
The last equality holds because for any $p\in P$, $\sum_{\tilde{s}\in S_{opt}}\phi_{S_{opt}}^{*}(p,\tilde{s})=1$ and $\sum_{\tilde{t}\in T}\phi_{T}^{*}(p,t)=1$. The first term is exactly $\rho$ times of (a) and the second term equals $\rho\cdot OPT$. Through combining (a) and (b), we can obtain an approximation factor of $\eta+(\eta+1)\rho$. ∎
$k$-Means in metric space.
Using the same idea of Lemma 7 with squared triangle inequality $\mathtt{dist}^{2}(a,b)\leq 2\mathtt{dist}^{2}(a,c)+2\mathtt{dist}^{2}(c,b)$, we can immediately obtain the following corollary.
Corollary 1.
Let $\eta$ be any positive number. If we suppose $\mathtt{Cost}(P,T,\phi_{T}^{*})\leq\eta\cdot OPT$, then the solution $(S,\phi_{S}^{*})$ returned by Algorithm 1 (slightly changed as above) is an $\big{(}2\eta+(4\eta+4)\rho\big{)}$-approximate solution for fair $k$-means problem in metric space, where $\rho$ is the approximation ratio of vanilla $k$-means.
When considering $k$-clustering problem in metric space, we usually assume that the potential facility set is given. We just use it as our candidate set $T$. Hence, the $\eta=1$ in the above analysis, which leads a $(2+\rho)$-approximation for fair $k$-median and a $(2+8\rho)$-approximation for fair $k$-means.
Appendix F Supplementary Experiment
F.1 Datasets
@@
F.2 Comparison on Cost with different $k$ and $(\alpha,\beta)$
@@
F.3 Comparison on Cost of $k$-sparse Wasserstein Barycenter
@@
F.4 Cost on Different sampling ratio
@@
F.5 Running time with different sampling ratio on $T$
As we discussed in F.4, sampling on relaxed solution $T$ can reduce the running time while the overall cost not increasing too much. We also test the running time with different sampling ratio. In summary, the running time of solving LP(2) on $T$ and overall Algorithm 1, shown in Table 3 and Table 4, can be significantly reduced by sampling.
@@ F.5 Running time with different sampling ratio on $T$
F.6 Comparison of Running Time with baselines
We compared the running time of our algorithm (Algorithm 1 with our rounding technique) with the baseline NIPS19 [Bera et al., 2019]. For strictly fair datasets, we also tested the running time of Algorithm 2 and ORL21 [Böhm et al., 2021]. The results are shown in Table 5 and Table 6. Below, we provide a detailed analysis on the comparisons.
Comparison between Algorithm 1 and NIPS19 [Bera et al., 2019].
Algorithm 1 and NIPS19 both have two important subprocedures: linear programming and the $k$-means algorithm. These two steps are the bottlenecks for Algorithm 1 and NIPS19. Specifically, NIPS19 first runs the $k$-means algorithm (i.e., $k$-means++), and then calls the LP solver once to compute the fractional assignment. A different part of our Algorithm 1 is that it calls the LP solver twice, once to compute the weights of candidate set $T$ and once to compute the fractional assignment, and calls the $k$-means algorithm once. In Algorithm 1, we only need to run $k$-means on $T$, which should be much smaller than the whole dataset, leading to less running time for the $k$-means subprocedure compared to NIPS19. However, the first call to the LP solver to compute the weight of $T$ consumes more time than the second call because $|T|>k$ usually. We illustrate the running time of every critical subprocedure of both algorithms in Table 5. Our $k$-means step is faster, but we have to run an extra LP step. Therefore, the running time comparison between these two algorithms is complex. Generally speaking, LP takes more time than $k$-means, which means our Algorithm 1 usually runs slower than NIPS19. However, with the development of LP solvers, we can expect that the runtime of Algorithm 1 could be further reduced with more advanced LP solvers.
Construct T LP on T k-means LP on S Rounding Total
Bank Algorithm1 0.01 2.4 $<$0.01 1.23 <0.01 3.78
NIPS19 / / 0.14 0.81 <0.01 1.11
Creditcard Algorithm 1 0.01 4.06 $<$0.01 2.27 <0.01 6.51
NIPS19 / / 0.18 2.05 <0.01 2.39
Census1990 Algorithm 1 0.01 7.51 0.02 5.19 <0.01 12.99
NIPS19 / / 0.30 3.94 <0.01 4.42
Adult Algorithm 1 0.01 4.14 $<$0.01 1.80 <0.01 6.12
NIPS19 / / 0.18 1.23 <0.01 1.59
Breastcancer Algorithm 1 0.01 0.19 $<$0.01 0.82 <0.01 1.33
NIPS19 / / 0.10 0.22 <0.01 0.45
Table 5: Running time (s) on non-strictly fair datasets
Table 6: Running time (s) on strictly fair datasets
Construct T LP on T k-means LP on S Rounding Total
Moons Algorithm 1 0.01 0.18 $<$0.01 0.64 <0.01 0.83
NIPS19 / / 0.07 0.70 0.01 0.78
Algorithm 2 / / $<$0.01 / / 0.59
ORL21 / / 0.02 / / 0.48
Cluto Algorithm 1 0.01 1.01 $<$0.01 1.30 <0.01 2.36
NIPS19 / / 0.07 1.54 <0.01 1.66
Algorithm 2 / / $<0.01$ / / 0.56
ORL 21 / / 0.56 / / 0.72
Complex Algorithm 1 0.01 1.08 $<$0.01 0.61 <0.01 1.71
NIPS19 / / 0.05 0.72 <0.01 0.79
Algorithm 2 / / $<0.01$ / / 0.58
ORL21 / / 0.56 / / 0.72
Hypercube Algorithm 1 0.01 5.71 0.01 4.40 <0.01 10.27
NIPS19 / / 0.15 2.58 <0.01 2.87
Algorithm 2 / / $<0.01$ / / 0.39
ORL21 / / 0.67 / / 0.83
Discussion on the construction of $T$.
According to Algorithm 1, $T$ should be an approximate centroid set [Matoušek, 2000]. Thanks to the open-source project by [Kanungo et al., 2002], which provides an efficient implementation of the approximate centroid set, we used their algorithm as part of our procedure in our code. Kanungo et al. [2002] used a sampling technique, leading to a trade-off between performance and efficiency. In our experiment, we sampled 10% of points in the approximate centroid set as $T$. A higher sample rate yields better performance (lower cost) but longer running time.
Besides, an implicit benefit of the construction of $T$ is that it is irrelevant to the parameters $k$, $\alpha$, and $\beta$. So if we consider a real scenario that we need to repeatedly try different choices for these parameters (e.g., we may want to tune the value $k$ and select the most satisfying result), the step of constructing $T$ and performing linear programming on $T$ can be seen as preprocessing of datasets before the tuning. Namely, we just need to run this preprocessing one time, and consequently the amortized cost over the whole tuning procedure can be reduced significantly.
Running time comparison on strictly fair datasets.
For strictly fair datasets, we consider Algorithm 1, NIPS19, Algorithm 2, and ORL21 . Algorithm 2 has an advantage in efficiency in most datasets. The primary reason is that Algorithm 2 only calls the $k$-means algorithm once and does not need to solve the LP. As for ORL21, it needs to run $k$-means for each group and then choose the best one. As a result, ORL21 takes longer time than Algorithm 2, especially on the datasets with large number of groups.
F.7 Experiments of Our Rounding Algorithm
@@
In this section, we implement our rounding algorithm in Appendix C and compute the violation factor across different datasets and parameters. For convenience, we parameterize $\alpha_{i}$ and $\beta_{i}$ for the $i$-th group using a single parameter $\delta$. Specifically, we set $\beta_{i}=\frac{|P^{(i)}|(1-\delta)}{|P|}$ and $\alpha_{i}=\frac{|P^{(i)}|}{|P|(1-\delta)}$. Generally speaking, the smaller the $\delta$, the stricter the fairness constrains are. In Table 7 8 9, the violation introduced by our rounding algorithm is less than 1 in most of the cases and never exceeds 2, which aligns with our theoretical analysis.